OpenStudy (anonymous):

the product of three intergers is 90. the second number is twice the first number. the third number is two more then the first number. what are the three numbers

OpenStudy (anonymous):

x(2x)(x+2)=90 can you solve this

OpenStudy (anonymous):

no

OpenStudy (anonymous):

2x^2(x+2)=90 2x^3+4x^2=90 2x^3+4x^2-90=0 how about now

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

no i cant

OpenStudy (anonymous):

can you help